# Redshift Dipoles from Non-Geodesic Observer Congruences in Covariant Cosmology

### 21 Mar 2026 v1

@@ Appendix A Evolution equation of $\sqrt{A}$
$\theta_{N}\equiv\frac{1}{A}\frac{dA}{d\lambda}.$ (A.1)
@@
$X(\lambda)\equiv\sqrt{A(\lambda)},\qquad\Rightarrow\qquad A=X^{2}.$ (A.2)
@@
$\frac{dA}{d\lambda}=2X\frac{dX}{d\lambda}.$ (A.3)
Substituting (A.3) into (A.1) yields
$\theta_{N}=\frac{1}{X^{2}}\left(2X\frac{dX}{d\lambda}\right)=2\,\frac{1}{X}\frac{dX}{d\lambda}.$ (A.4)
Now differentiate (A.4) once more:
@@
$=2\left[\frac{1}{X}\frac{d^{2}X}{d\lambda^{2}}-\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}\right].$ (A.5)
On the other hand, squaring (A.4) gives
$\frac{1}{2}\,\theta_{N}^{2}=\frac{1}{2}\left(4\,\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}\right)=2\,\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}.$ (A.6)
Adding (A.5) and (A.6), the quadratic terms cancel identically, leaving
$\frac{d\theta_{N}}{d\lambda}+\frac{1}{2}\,\theta_{N}^{2}=2\,\frac{1}{X}\frac{d^{2}X}{d\lambda^{2}}.$ (A.7)
@@
$\boxed{\frac{d\theta_{N}}{d\lambda}+\frac{1}{2}\,\theta_{N}^{2}=\frac{2}{\sqrt{A}}\frac{d^{2}\sqrt{A}}{d\lambda^{2}}.}$ (A.8)
@@ Appendix B Proof that $d_{A}=aS_{K}(\chi)$ solves the Sachs equation in FLRW
$d\eta^{2}+d\chi^{2}=0\qquad\Rightarrow\qquad\frac{d\chi}{d\eta}=\pm 1.$ (B.1)
@@
Using (B.2), we obtain
$k^{\chi}=\frac{d\chi}{d\eta}\frac{d\eta}{d\lambda}=-\,k^{\eta}.$ (B.3)
@@
$\frac{df}{d\lambda}=k^{\alpha}\nabla_{\alpha}f=k^{\eta}\partial_{\eta}f+k^{\chi}\partial_{\chi}f.$ (B.4)
Substituting (B.3), we find
$\frac{df}{d\lambda}=k^{\eta}\partial_{\eta}f-k^{\eta}\partial_{\chi}f=k^{\eta}\left(\partial_{\eta}-\partial_{\chi}\right)f.$ (B.5)
@@
$\frac{dd_{A}}{d\lambda}=k^{\eta}\left(\frac{da}{d\eta}\,S_{K}-a\,\frac{dS_{K}}{d\chi}\right).$ (B.6)
@@
$=\left(\frac{dk^{\eta}}{d\lambda}\right)\left(\frac{da}{d\eta}S_{K}-a\frac{dS_{K}}{d\chi}\right)+(k^{\eta})^{2}\left(\frac{d^{2}a}{d\eta^{2}}S_{K}-2\frac{da}{d\eta}\frac{dS_{K}}{d\chi}+a\frac{d^{2}S_{K}}{d\chi^{2}}\right).$ (B.7)
@@
$\frac{dk^{\eta}}{d\lambda}+\frac{2}{a}\frac{da}{d\eta}(k^{\eta})^{2}=0,\qquad\Rightarrow\qquad\frac{dk^{\eta}}{d\lambda}=-\frac{2}{a}\frac{da}{d\eta}(k^{\eta})^{2}.$ (B.8)
Substituting into (B.7), the terms proportional to $\frac{dS_{K}}{d\chi}$ cancel identically, yielding
$\frac{d^{2}d_{A}}{d\lambda^{2}}=(k^{\eta})^{2}\left[a\frac{d^{2}S_{K}}{d\chi^{2}}+\left(\frac{d^{2}a}{d\eta^{2}}-2\frac{(da/d\eta)^{2}}{a}\right)S_{K}\right].$ (B.9)
@@
$\frac{d^{2}d_{A}}{d\lambda^{2}}=(k^{\eta})^{2}S_{K}\left[\frac{d^{2}a}{d\eta^{2}}-\frac{2}{a}\left(\frac{da}{d\eta}\right)^{2}-aK\right]=-(k^{\eta})^{2}\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}-\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}+K\right]d_{A}.$ (B.11)
@@
$\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}+K\right]=\frac{8\pi G}{3}a^{2}\rho+\frac{\Lambda}{3}a^{2},$ (B.12)
@@
$\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}=-\frac{4\pi G}{3}a^{2}(\rho+3p)+\frac{\Lambda}{3}a^{2}.$ (B.13)
Combining (B.12)–(B.13) gives the useful identity
$\boxed{\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}-\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}+K\right]=4\pi Ga^{2}(\rho+p).}$ (B.14)
Substituting Eq. (B.14) into Eq. (B.11) we obtain
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-4\pi G\,a^{2}(\rho+p)\,(k^{\eta})^{2}\,d_{A}.$ (B.15)
@@
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)\,a^{2}\,(k^{\eta})^{2}.$ (B.16)
Comparing this with Eq. (B.15) immediately gives
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-\frac{1}{2}\,R_{\alpha\beta}k^{\alpha}k^{\beta}\,d_{A},$ (B.17)
@@ C.1 Homogeneous scalars in FLRW
$h_{\alpha}{}^{\beta}\nabla_{\beta}a=0,$ (C.1)
@@
$\nabla_{\alpha}a=-u_{\alpha}\dot{a}+h_{\alpha}{}^{\beta}\nabla_{\beta}a,\qquad\dot{a}\equiv u^{\beta}\nabla_{\beta}a,$ (C.2)
@@
$\nabla_{\alpha}a=-\dot{a}\,u_{\alpha}.$ (C.3)
@@ C.2 Contractions involving the photon wavevector
$k^{\alpha}=E(u^{\alpha}+e^{\alpha}),\qquad u_{\alpha}e^{\alpha}=0,\qquad e_{\alpha}e^{\alpha}=1,\qquad u_{\alpha}u^{\alpha}=-1.$ (C.4)
@@
$k^{\alpha}k^{\beta}h_{\alpha\beta}=E^{2}(u^{\alpha}+e^{\alpha})(u^{\beta}+e^{\beta})h_{\alpha\beta}=E^{2}.$ (C.5)
@@
$k^{\alpha}k^{\beta}\sigma_{\alpha\beta}=E^{2}\sigma_{\alpha\beta}e^{\alpha}e^{\beta}.$ (C.6)
@@
$k^{\alpha}k^{\beta}\omega_{\alpha\beta}=0.$ (C.7)
@@
$k^{\alpha}k^{\beta}u_{\alpha}\mathcal{A}_{\beta}=E^{2}(u^{\alpha}+e^{\alpha})(u^{\beta}+e^{\beta})u_{\alpha}\mathcal{A}_{\beta}=-E^{2}\,\mathcal{A}_{\alpha}e^{\alpha}.$ (C.8)
@@ C.3 Relation between affine parameter and observer time
$E\equiv-(u_{\alpha}k^{\alpha}).$ (C.9)
@@
$k^{\alpha}=\frac{dx^{\alpha}}{d\lambda}=\left(\frac{dt}{d\lambda},\frac{dx^{i}}{d\lambda}\right).$ (C.11)
@@
$E=-u_{\alpha}k^{\alpha}=\frac{dt}{d\lambda},$ (C.12)
@@
$\frac{d}{d\lambda}=\frac{dt}{d\lambda}\frac{d}{dt}=E\,\frac{d}{dt}.$ (C.13)

### 14 Apr 2026 v2

@@ Appendix A Evolution equation of $\sqrt{A}$
$\theta_{N}\equiv\frac{1}{A}\frac{dA}{d\lambda}.$ (63)
@@
$X(\lambda)\equiv\sqrt{A(\lambda)},\qquad\Rightarrow\qquad A=X^{2}.$ (64)
@@
$\frac{dA}{d\lambda}=2X\frac{dX}{d\lambda}.$ (65)
Substituting (65) into (63) yields
$\theta_{N}=\frac{1}{X^{2}}\left(2X\frac{dX}{d\lambda}\right)=2\,\frac{1}{X}\frac{dX}{d\lambda}.$ (66)
Now differentiate (66) once more:
@@
$=2\left[\frac{1}{X}\frac{d^{2}X}{d\lambda^{2}}-\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}\right].$ (67)
On the other hand, squaring (66) gives
$\frac{1}{2}\,\theta_{N}^{2}=\frac{1}{2}\left(4\,\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}\right)=2\,\frac{1}{X^{2}}\left(\frac{dX}{d\lambda}\right)^{2}.$ (68)
Adding (67) and (68), the quadratic terms cancel identically, leaving
$\frac{d\theta_{N}}{d\lambda}+\frac{1}{2}\,\theta_{N}^{2}=2\,\frac{1}{X}\frac{d^{2}X}{d\lambda^{2}}.$ (69)
@@
$\boxed{\frac{d\theta_{N}}{d\lambda}+\frac{1}{2}\,\theta_{N}^{2}=\frac{2}{\sqrt{A}}\frac{d^{2}\sqrt{A}}{d\lambda^{2}}.}$ (70)
@@ Appendix B Proof that $d_{A}=aS_{K}(\chi)$ solves the Sachs equation in FLRW
$d\eta^{2}+d\chi^{2}=0\qquad\Rightarrow\qquad\frac{d\chi}{d\eta}=\pm 1.$ (71)
@@
Using (72), we obtain
$k^{\chi}=\frac{d\chi}{d\eta}\frac{d\eta}{d\lambda}=-\,k^{\eta}.$ (73)
@@
$\frac{df}{d\lambda}=k^{\alpha}\nabla_{\alpha}f=k^{\eta}\partial_{\eta}f+k^{\chi}\partial_{\chi}f.$ (74)
Substituting (73), we find
$\frac{df}{d\lambda}=k^{\eta}\partial_{\eta}f-k^{\eta}\partial_{\chi}f=k^{\eta}\left(\partial_{\eta}-\partial_{\chi}\right)f.$ (75)
@@
$\frac{dd_{A}}{d\lambda}=k^{\eta}\left(\frac{da}{d\eta}\,S_{K}-a\,\frac{dS_{K}}{d\chi}\right).$ (76)
@@
$=\left(\frac{dk^{\eta}}{d\lambda}\right)\left(\frac{da}{d\eta}S_{K}-a\frac{dS_{K}}{d\chi}\right)+(k^{\eta})^{2}\left(\frac{d^{2}a}{d\eta^{2}}S_{K}-2\frac{da}{d\eta}\frac{dS_{K}}{d\chi}+a\frac{d^{2}S_{K}}{d\chi^{2}}\right).$ (77)
@@
$\frac{dk^{\eta}}{d\lambda}+\frac{2}{a}\frac{da}{d\eta}(k^{\eta})^{2}=0,\qquad\Rightarrow\qquad\frac{dk^{\eta}}{d\lambda}=-\frac{2}{a}\frac{da}{d\eta}(k^{\eta})^{2}.$ (78)
Substituting into (77), the terms proportional to $\frac{dS_{K}}{d\chi}$ cancel identically, yielding
$\frac{d^{2}d_{A}}{d\lambda^{2}}=(k^{\eta})^{2}\left[a\frac{d^{2}S_{K}}{d\chi^{2}}+\left(\frac{d^{2}a}{d\eta^{2}}-2\frac{(da/d\eta)^{2}}{a}\right)S_{K}\right].$ (79)
@@
$\frac{d^{2}d_{A}}{d\lambda^{2}}=(k^{\eta})^{2}S_{K}\left[\frac{d^{2}a}{d\eta^{2}}-\frac{2}{a}\left(\frac{da}{d\eta}\right)^{2}-aK\right]=-(k^{\eta})^{2}\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}-\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}+K\right]d_{A}.$ (81)
@@
$\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}+K\right]=\frac{8\pi G}{3}a^{2}\rho+\frac{\Lambda}{3}a^{2},$ (82)
@@
$\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}=-\frac{4\pi G}{3}a^{2}(\rho+3p)+\frac{\Lambda}{3}a^{2}.$ (83)
Combining (82)–(83) gives the useful identity
$\boxed{\left[\left(\frac{1}{a}\frac{da}{d\eta}\right)^{2}-\frac{1}{a}\frac{d^{2}a}{d\eta^{2}}+K\right]=4\pi Ga^{2}(\rho+p).}$ (84)
Substituting Eq. (84) into Eq. (81) we obtain
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-4\pi G\,a^{2}(\rho+p)\,(k^{\eta})^{2}\,d_{A}.$ (85)
@@
$R_{\alpha\beta}k^{\alpha}k^{\beta}=8\pi G(\rho+p)\,a^{2}\,(k^{\eta})^{2}.$ (86)
Comparing this with Eq. (85) immediately gives
$\frac{d^{2}d_{A}}{d\lambda^{2}}=-\frac{1}{2}\,R_{\alpha\beta}k^{\alpha}k^{\beta}\,d_{A},$ (87)
@@ C.1 Homogeneous scalars in FLRW
$h_{\alpha}{}^{\beta}\nabla_{\beta}a=0,$ (88)
@@
$\nabla_{\alpha}a=-u_{\alpha}\dot{a}+h_{\alpha}{}^{\beta}\nabla_{\beta}a,\qquad\dot{a}\equiv u^{\beta}\nabla_{\beta}a,$ (89)
@@
$\nabla_{\alpha}a=-\dot{a}\,u_{\alpha}.$ (90)
@@ C.2 Contractions involving the photon wavevector
$k^{\alpha}=E(u^{\alpha}+e^{\alpha}),\qquad u_{\alpha}e^{\alpha}=0,\qquad e_{\alpha}e^{\alpha}=1,\qquad u_{\alpha}u^{\alpha}=-1.$ (91)
@@
$k^{\alpha}k^{\beta}h_{\alpha\beta}=E^{2}(u^{\alpha}+e^{\alpha})(u^{\beta}+e^{\beta})h_{\alpha\beta}=E^{2}.$ (92)
@@
$k^{\alpha}k^{\beta}\sigma_{\alpha\beta}=E^{2}\sigma_{\alpha\beta}e^{\alpha}e^{\beta}.$ (93)
@@
$k^{\alpha}k^{\beta}\omega_{\alpha\beta}=0.$ (94)
@@
$k^{\alpha}k^{\beta}u_{\alpha}\mathcal{A}_{\beta}=E^{2}(u^{\alpha}+e^{\alpha})(u^{\beta}+e^{\beta})u_{\alpha}\mathcal{A}_{\beta}=-E^{2}\,\mathcal{A}_{\alpha}e^{\alpha}.$ (95)
@@ C.3 Relation between affine parameter and observer time
$E\equiv-(u_{\alpha}k^{\alpha}).$ (96)
@@
$k^{\alpha}=\frac{dx^{\alpha}}{d\lambda}=\left(\frac{dt}{d\lambda},\frac{dx^{i}}{d\lambda}\right).$ (98)
@@
$E=-u_{\alpha}k^{\alpha}=\frac{dt}{d\lambda},$ (99)
@@
$\frac{d}{d\lambda}=\frac{dt}{d\lambda}\frac{d}{dt}=E\,\frac{d}{dt}.$ (100)